Eugene spent $14 on a magazine and some candy bars.If the magazine cost $6 and each candy bar cost $2 then how much candy bars did he buy?
I tried setting it up by
6+2x=14
-6. -6
2x=8
2. 2
X=4

Found 3 solutions by Shin123, amalm06, ikleyn:
Answer by Shin123(626)

6+2x=14
-6=0 -6=8 2x =8
x =4

Answer by amalm06(224)

This is an optimization problem.
Budget constraint=$14.
$6 is spent on a magazine.
$14-$6=$8 is leftover.
If one candy bar is $2, what is the maximum number of candy bars that can be purchased with $8? The answer is 4.

Algebraically, 6+2x=14
2x=8
x=4 candy bars

A second method of solution is that of alligation.
Denote magazines by M and candy bars by C.
Then 14-2=12 and 14-6=8
So that M/C=12/8=3/2 or C=(2/3)M
E.g. the number of candy bars is 2/3 the price of a magazine
Since one magazine costs $6, C=(2/3)(6)=4 candy bars

Answer by ikleyn(52756)

.
Your steps, your solution and your answer are all absolutely correct.


1.  You correctly wrote the equation 6 + 2x = 14.


2.  You correctly subtracted 6 from both sides and obtained the equivalent equation 2x = 8.


3.  You correctly divided both sides by 2 and obtained the solution x = 4.


CONGRATULATIONS !!

If you want, you may look into the lessons
    - HOW TO solve a linear equation
    - Using linear equations to solve word problems
    - More solved word problems on a single linear equation
    - OVERVIEW of lessons on solving linear equations and word problems in one unknown
in this site.

These lessons are specially intended for those students who only start learning equations and their solutions.


                    H a p p y   l e a r n i n g  ! !


---------------

Look into the solution of the young tutor  @Shin123  !

Her ideas and the way of thinking are very relevant to the given problem.


Unfortunately,  I can not say the same about writing by  @amalm06.